# Assessing the quality of information extraction

### v1

@@ Abstract
Advances in large language models have notably enhanced the efficiency of information extraction from unstructured and semi-structured data sources. As these technologies become integral to various applications, establishing an objective measure for the quality of information extraction becomes imperative. However, the scarcity of labeled data presents significant challenges to this endeavor. In this paper, we introduce an automatic framework to assess the quality of the information extraction and its completeness. The framework focuses on information extraction in the form of entity and its properties. We discuss how to handle the input/output size limitations of the large language models and analyze their performance when iteratively extracting the information. Finally, we introduce metrics to evaluate the quality of the extraction and provide an extensive discussion on how to interpret the metrics.
Keywords: information extraction, large language models, quality evaluation, name entity recognition, needle in a haystack test, schema.org
@@ 1 Introduction
In the domain of natural language processing (NLP), information extraction (IE) stands as a critical task, transforming unstructured or semi-structured data into a structured format conducive to indexing, exploration, and further analysis. The increasing amount of data across digital platforms underscores the urgency for sophisticated IE techniques that can parse through volumes of information with precision. An extensive survey about IE is provided by, where the authors highlight the complexity of processing and analyzing text to derive meaningful information, given the heterogeneity and volume of such data.
Large language models (LLMs) have revolutionized IE by introducing generative methods for structuring knowledge from text. LLMs excel across diverse domains without extensive task-specific training. A survey by [8] details the progress of LLMs on IE tasks. Here, the authors address specific aspects of information extraction, including entity recognition, relation extraction, event detection, and universal IE. They review the existing models and their efficiency on a comprehensive collection of annotated benchmarks. Nonetheless, the challenge of quantitatively assessing the quality and completeness of extracted information persists, particularly in the absence of labeled datasets for benchmarking. Before conducting the experiments introduced in this paper, we perform IE on a vast corpus of business documents utilizing LLMs. While the extraction process is beyond the scope of this paper, some detail about the extraction is given in Section 3.
To measure the quality of extraction, we propose an evaluation framework that relies on artificially generated complex information which is infused into the document to test the efficiency of LLMs in IE tasks. This paper introduces an iterative extraction process and a novel score, MINEA (Multiple Infused Needle Extraction Accuracy), to address the critical need for objective quality assessment measures. By inserting artificial information (”needles”) into the data, the proposed method creates a synthetic ground truth for evaluation, enabling the measurement of extraction quality in various specific domains even without manually labeled data. The empirical analysis demonstrates the utility of MINEA for evaluating LLM-based IE in scenarios where ground truth is unavailable.
This paper begins a trilogy on our GRIX framework, focusing on evaluation methods for our upcoming discussions on advanced extraction, indexing, and retrieval techniques.
@@ 2 Related work
A common practice in many specialized IE tasks is that well-trained experts review what was extracted, and provide ground truth [4]. Such an approach is relatively reliable, however, it is manual and very time-consuming.
In [3] they suggest summary score without reference (SUSWIR), a score to evaluate the quality of text summaries without the need for human annotations. The SUSWIR score can be used for IE tasks where the extracted information can be viewed as a compression of initial data. The score compares the original text with its summary. From its nature, it is very useful when comparing the outputs of extraction among themselves, i.e., the best extraction/summary has the highest score value. On the other hand, its ability to provide an objective absolute evaluation of a single extraction is disadvantaged because the desirable output is not known.
@@ 3 Capturing the structure
In this section, we explain how structured information can be obtained from a text document by an LLM.
@@ 3.1 Schema
To impose a structure on the data, we adopt the idea of schema markup [2] which is used to communicate the content of a web page to the search tool. The schema markup is in the form of structured data and can be viewed as a compression of the essential information. The structure is defined by Schema.org [^3] https://schema.org vocabulary which is a set of entity types, each associated with a set of properties and hierarchically arranged. Figure 1 shows an example of structured information. It describes three entities of types ‘Insight’, ‘Person’ and ‘Organization’. Each type has its own set of properties, e.g., an entity of type ‘Person’ is described by ‘type’, ‘name’, ‘birthDate’, ‘worksFor’, and ‘jobTitle’. In other words, each entity is a set of key-value pairs, e.g., ‘name’ is the key and ‘AI Enthusiast’ is the value.
@@
Similarly, we extract and compress the relevant information contained in the data using an LLM. Schema.org presents a clear basis for the categorization of various entities contained in the data. In the rest of the paper, by schema we mean a predetermined set of types, such as {‘Person’, ‘Project’, ‘Product’, ‘Legislation’, ‘Event’, ‘OpportunityArea’, ‘Insight’, ‘Substance’, ‘Thing’, ‘BioChemEntity’, ‘MedicalCondition’}. The schema is set at the beginning and the information to be extracted depends on it. Therefore the schema has to be tailored to a particular scope of the (proprietary) knowledge and application. If a more complex or uncommon entity needs to be captured, it is natural and very easy to extend the set of core types by more detailed descriptive and custom vocabulary. E.g., ‘Insight’ and ‘OpportunityArea’ are not native schema.org types, but we will use them in our study. The usage of suitably tailored schema is beneficial for specialized applications since it narrows the information to the relevant core and hence potentially improves the overall performance. On the other hand, the usage of schemata is not restrictive as the scope can be always extended by using a broader set of types.
@@ 3.2 The role of LLMs
LLMs are rather effective with the creation of structured data with predefined types and attributes (properties), cf. [8]. Together with dedicated prompts, we get a structured file describing entities found in the documents and matching schema types. Besides the extraction task, LLMs can be used to suggest suitable schema.org types for a particular document. An example together with the prompt is shown in Appendix B1.
@@ 4 Length aspects
When focusing on the quality of IE using LLMs we need to take into account several limitations posed by LLMs to the length of the data to extract from. Each LLM has a maximal content limit it can process, both on the input and the output. The limit on the output is typically much more strict. When trying to use the maximal possible input another issue may appear – the Lost in the middle phenomenon [7] says that the ability of LLMs to retrieve information from a long context declines and that the attention focuses on the beginning and the end of the context while it tends to attenuate information in the middle.
To demonstrate this shortcoming numerically we use gpt-4-1106-preview model [^4] https://platform.openai.com/docs/models/overview. The model is limited by 4095 tokens on the output and by 128000 tokens on the input (context window limit). Further, note that OpenAI model versions $>4$ allow outputting the response directly in a structured format.
The following sections present two major LLM limitations we have to consider before performing IE, namely length restrictions in Section 4.1 and Lost in the middle problem in Section 4.2.
@@ 4.1 Length restrictions
    output tokens: The restriction on output tokens means that there is some maximum length of data to effectively extract the majority of entities from it. If the length of the text exceeds this maximum, there would be no tokens for extra entities.
@@
Another difficulty regarding the output is the tendency of LLMs to generate rather brief responses which do not use the allowed maximum number of tokens. This unwillingness of models can be circumvented by prompting. Even so, the limited number of output tokens is typically too low and prevents effective extraction from long texts.
With a more sophisticated approach, the restriction (O) becomes irrelevant and only the restriction (I) will apply. The issue imposed by (O) is overcome by splitting the source document into smaller pieces which are extracted independently. A significant drawback is that the extracted information can be easily duplicated – extracted independently from multiple text pieces. Iterating the calls to the LLM with instruction to continue with already started extraction, i.e., continuing with the extraction in a single thread, helps to extract more information and to avoid duplication. Since we insist on continuation more and more information is added and the extraction is more thorough, at least to some point – this will be addressed in detail in Section 5.1. Further, a lower number of duplicates is found due to the extraction history, i.e., all information extracted until present, which is kept within the thread.
@@ 5 Quality of extraction
Once the information is extracted from an initial text document into a structured form defined by the chosen schema, e.g., Figure 1, the quality of such extraction is important to evaluate. In practice, it is very rare to be equipped with ground truth and its human generation requires vast expertise in the scope of the document and a ridiculous amount of time. Therefore we adopt methods from [3]. They examine semantic similarity, relevance, redundancy, and bias and compound these into a single score called SUSWIR, for details see Appendix A. The score and its subparts are very useful when comparing distinct extractions among themselves, e.g., we can use it to find an optimal number of iterated LLM calls. Unfortunately, the score does not represent an absolute way of evaluation. It does not provide a complete insight into the task – some information = entities can be missing, misclassified or their properties not filled in correctly. To come up with a robust and general solution we generalize the NIAH test, which is commonly used to measure the ability of LLMs to process long documents, cf. [5].
@@ 5.1 Iterated LLM calls
Since the first LLM extraction is typically not complete, iterating the extraction process helps with the completeness of extraction. To improve the quality of extraction, we ask LLM to process the document again and search for other entities which were not extracted yet. A question arises: What is the optimal number of iterations? It is desirable to stop when additional LLM call will return no or only a few new entities. The answer however depends heavily on the text being extracted and on the chosen schema.
Below, we present a small comparative study regarding the contribution of iterated extraction to its quality. We interpret the extracted structured data, e.g., Figure 3, as a summary of the original text document. To measure the quality of the summary we adopt the scores from [3] (a convex combination of these scores creates the overall SUSWIR metric), namely semantic similarity, relevance, and redundancy avoidance. We use a modified bias avoidance score from [3] and add two new scores, relevance spread, and incompleteness score.
Table 2 compares the initial text document with extracted information created iteratively by succeeding LLM calls. Each iteration enriches the extracted information, but the benefit decreases. From the third iteration, i.e., after four LLM calls, the majority of scores in Table 2 are either getting worse or stagnating (the arrows following the score name indicate the direction in which the score improves).
@@
The main conclusion of Table 2 is that iterating the LLM calls has limits. From some point, the extracted information is more complete and thorough neither in a semantic nor factually relevant sense. Further, the risk that the LLM will suffer from hallucinations increases as we observe a growth of bias.
@@ 5.2 Test the quality
This section introduces a robust and versatile score to objectively measure the quality of IE. Assuming the structure is imposed by some schema, see Section 3.1, we would like to measure the IE quality as a portion of successfully extracted entities, i.e., the accuracy of name entity recognition (NER) task taking into account even the context captured by entity properties. Unfortunately, such an experiment is unfeasible without labeled data. As a consequence, it is unfeasible in many specialized tasks because of the absence of suitable labeled data unseen by LLM models. This can be the case witf very recent datasets as well as proprietary datasets. To overcome this issue we use inspiration from NIAH test to build up an automatic and general procedure to access the quality of IE.
@@ 5.2.1 Needles
The ‘needle’ in our context represents the entity. It is created according to the chosen schema, i.e., a list of types we want to extract from the document. We use an LLM to generate a short paragraph introducing a new original (not appearing in the document) entity, but still relevant to the scope of the document, for an example see Figure 2, and for more details see Appendix B2. This artificial paragraph, the needle, is then placed into the document body at random (taking into the account natural units within the text as sentences, paragraphs, etc. if applicable). Moreover, the needle is accompanied with several properties, namely we assign to the needle a name, short description and keywords, see Figure 2. This additional properties are assigned to the needle by the LLM.
@@
5.2.2 Multiple needle retrieval accuracy
To measure the quality of extraction we propose a multiple infused needle extraction accuracy (MINEA) score. Its computation combines the approach of NIAH evaluation and NER task. We scatter several needles over the text document body and measure how many of them were successfully extracted. Since we know what exactly was inserted, we know what should be extracted. Then we can objectively measure the quality of extraction on these new entities and moreover, we can compare extracted information from the document with and without needles. Table 3 shows extraction accuracy – MINEA score – total and per schema type – measured on a vast corpus of business documents with predefined schema consisting of types ‘BioChemEntity’, ‘Event’, ‘Insight’, ‘Legislation’, ‘MedicalCondition’, ‘OpportunityArea’, ‘Person’, ‘Product’, ‘Project’, ‘Substance’ and ‘Thing’.
@@ 5.2.3 Identification of needles
Matching the generated needles with extracted entities imposes a challenge and mostly depends on the formulation of needles. If the needles are too complex or too vague the straightforward identification changes into a serious problem. For this reason, we equip the needles with additional properties which are then used to compare the needles with extracted entities and to decide whether the needles were extracted or not.
@@
    an entity with some number of keywords perfectly matching the needle keywords is found, the number is determined by the threshold parameter determining the percentage amount of keywords to be matched;
@@
Note that other conditions can be constructed, e.g., based on the short description instead of keywords. Table 4 shows whether the conditions are fulfilled in the example illustrated by Figures 2 and 3. Namely, the condition n is not satisfied (‘AI Clan Meeting’ $\neq$ ‘AI Meeting’, ‘Graph Index’ $\neq$ ‘GRIX’). Condition ns is satisfied only for needle representing an entity of type ‘Event’ (‘AI Clan Meeting’ can be found in the extracted information). There are three keywords from those six assigned to the needle representing the entity of type ‘Event’ which match the keywords of an extracted entity, hence k0.5 is, and k0.6, k0.7 are not satisfied (there is an entity within the extracted information with $50\%$ of keywords being the same as the keywords of the needle). In the case of the second needle, there are four such keywords, therefore k0.5 and k0.6 are satisfied. Finally, both needles are identified within the extracted information by an LLM.
Table 5 shows scores (ratios of successfully extracted entities) based on the above criteria in the case of our business documents. The highest score per entity type is highlighted. Matching the needle and entity name usually does not perform well if the name is prone to modification (e.g., person name with and without title), or if the entity is easy to be misclassified (an entity of type ‘country’ was usually extracted as ‘place’ whose name did not match the country name). Searching for the needle name within the all extracted information gives very accurate results in case of entities being characterized well by their name (compare for example types ‘person’ and ’legislation’ with type ’insight’ where the name is not a natural attribute). Matching the needle and entity keywords depends on the threshold parameter – with a lower proportion of keywords that have to match the score value increases and the reliability of the entity identification decreases. An LLM performs well the entity identification and it is an important criterion in the case of more creative types such as ‘insight’. Finally, the MINEA score for each type is taken as the maximum of the scores.

### v2

@@ Abstract
Advances in large language models have notably enhanced the efficiency of information extraction from unstructured and semi-structured data sources. As these technologies become integral to various applications, establishing an objective measure for the quality of information extraction becomes imperative. However, the scarcity of labeled data presents significant challenges to this endeavor. In this paper, we introduce an automatic framework to assess the quality of the information extraction/retrieval and its completeness. The framework focuses on information extraction in the form of entity and its properties. We discuss how to handle the input/output size limitations of the large language models and analyze their performance when extracting the information. In particular, we introduce scores to evaluate the quality of the extraction and provide an extensive discussion on how to interpret them.
@@ 1 Introduction
In the domain of natural language processing (NLP), information extraction (IE) stands as a critical task, transforming unstructured or semi-structured data into a structured format conducive to indexing, exploration, and further analysis. The increasing amount of data across digital platforms underscores the urgency for sophisticated IE techniques that can parse through volumes of information with precision. An extensive survey about IE is provided by [1], where the authors highlight the complexity of processing and analyzing text to derive meaningful information, given the heterogeneity and volume of such data.
Large language models (LLMs) have revolutionized IE by introducing generative methods for structuring knowledge from text. LLMs excel across diverse domains without extensive task-specific training. A survey by [9] details the progress of LLMs on IE tasks. Here, the authors address specific aspects of information extraction, including entity recognition, relation extraction, event detection, and universal IE. They review the existing models and their efficiency on a comprehensive collection of annotated benchmarks. Nonetheless, the challenge of quantitatively assessing the quality and completeness of extracted information persists, particularly in the absence of labeled datasets for benchmarking. Before conducting the experiments introduced in this paper, we perform IE on a vast corpus of business documents utilizing LLMs. While the extraction process is beyond the scope of this paper, some details about the extraction are given in Section 3.
To measure the quality of extraction, we propose an evaluation framework that relies on artificially generated complex information which is infused into the document to test the efficiency of LLMs in IE tasks. This paper introduces an iterative extraction process and a novel score, MINEA (Multiple Infused Needle Extraction Accuracy), to address the critical need for objective quality assessment measures. By inserting artificial information ("needles") into the data, the proposed method creates a synthetic ground truth for evaluation, enabling the measurement of extraction quality in various specific domains even without manually labeled data. The empirical analysis demonstrates the utility of MINEA for evaluating LLM-based IE in scenarios where ground truth is unavailable. By automating the quality assessment of information extraction, the framework could reduce the need for manual review by experts, saving time and resources and thus enhance the efficiency and accuracy of information extraction from large volumes of unstructured data.
@@ 2 Related work
A common practice in many specialized IE tasks is that well-trained experts review what was extracted and provide ground truth as done in [5]. Such an approach is relatively reliable, however, it is manual and very time-consuming.
In [4] they suggest summary score without reference (SUSWIR), a score to evaluate the quality of text summaries without the need for human annotations. The SUSWIR score can be used for IE tasks where the extracted information is viewed as a compression of original data. The score compares the original text with its summary. From its nature, it is very useful when comparing the outputs of extraction tasks among themselves, i.e., the best extraction/summary has the highest score value. On the other hand, its ability to provide an objective absolute evaluation of a single extraction is disadvantaged because the desirable output is not known.
@@ 3 Capturing the structure
The form of needles depends on a form of data, on structure capturing the information and on the task being solved. The needles can be short paragraphs of text, account records, graph nodes as you extract information from continuous text, table, graph, respectively. The structured arrangement of information is beneficial for consecutive processing and analysis. It helps to highlight relationships among distinct information pieces. There are countless ways to impose a structure on unstructured data in order to capture the relevant information. To demonstrate our methodology for measuring the quality of information extraction, we specify a particular structure and tailor the needles to it.
@@ 3.1 Schema
To impose a structure on the data, we adopt the idea of schema markup [3] which is used to communicate the content of a web page to the search tool. The schema markup is in the form of structured data and can be viewed as a compression of the essential information. The structure is defined by Schema.org [^2] https://schema.org vocabulary which is a set of entity types, each associated with a set of properties and hierarchically arranged. Figure 1 shows an example of structured information inspired by Schema.org. It describes three entities of types ‘Insight’, ‘Person’ and ‘Organization’. Each type has its own set of properties, e.g., an entity of type ‘Person’ is described by ‘type’, ‘name’, ‘birthDate’, ‘worksFor’, and ‘jobTitle’. In other words, each entity is a set of key-value pairs, e.g., ‘name’ is the key and ‘AI Enthusiast’ is the value.
@@
Similarly, we extract and compress the relevant information contained in data using an LLM. Schema.org presents a clear basis for the categorization of various entities contained in data. In the rest of the paper, by schema we mean a predetermined set of types, such as {‘Person’, ‘Project’, ‘Product’, ‘Legislation’, ‘Event’, ‘OpportunityArea’, ‘Insight’, ‘Substance’, ‘Thing’, ‘BioChemEntity’, ‘MedicalCondition’}, together with their properties. The schema is set at the beginning and the information to be extracted depends on it. Therefore the schema has to be tailored to a particular scope of the (proprietary) knowledge and application. If a more complex or uncommon entity needs to be captured, it is natural and very easy to extend the set of core types by more detailed descriptive and custom vocabulary. E.g., ‘Insight’ and ‘OpportunityArea’ are not native Schema.org types, but we will use them in our study. The usage of suitably tailored schema is beneficial for specialized applications since it narrows the information to the relevant core and hence potentially improves the overall performance. On the other hand, the usage of schemata is not restrictive as the scope can be always extended by using a broader set of types.
@@ 3.2 The role of LLMs
LLMs are rather effective in the creation of structured data, cf. [9]. Using dedicated prompts, we get a structured text file describing entities found in the documents and matching types of predefined schema. The predefined schema (types and properties) is given to an LLM within the prompt. The LLM is asked to analyse the document, identify an information relevant to the mentioned types of entities and populate the schema with this information. It is asked to be attentive to nested entities, maintain consistency and uniqueness of extracted entities. Indeed, LLM is not prohibited from extracting entities whose types do not appear in the predefined schema. It is worthy to note, that LLMs are known to inherit biases present in their training data. If not carefully managed, these biases could lead to unfair or inaccurate information extraction, impacting decision-making processes.
Besides the information extraction task, LLMs can be used to suggest suitable Schema.org types for a particular document. An example together with a prompt is shown in Appendix B1.
@@ 4 Length aspects
When focusing on the quality of IE performed by an LLM, several limitations that LLM presents in terms of the length of data to be extracted from must be considered. Each LLM has a maximal content limit it can process, both on the input and the output. The limit on the output is typically much more strict. When trying to use the maximal possible input another issue may appear – the Lost in the middle phenomenon [8] says that the ability of LLMs to retrieve information from a long context declines and that the attention focuses on the beginning and the end of the context while it tends to attenuate information in the middle.
To demonstrate shortcomings arising from these limitations numerically we use gpt-4-1106-preview model. [^3] https://platform.openai.com/docs/models/overview The model is limited by 4095 tokens on the output and by 128000 tokens on the input (context window limit). The following sections present two major LLM limitations we have to consider before performing IE, namely length restrictions in Section 4.1 and Lost in the middle problem in Section 4.2.
@@ 4.1 Length restrictions
    output tokens: The restriction on output tokens means that there is some maximal length of data from which most entities can be efficiently extracted. If the length of the text exceeds this maximum, there would be no tokens for extra entities.
@@
Another difficulty regarding the output is the tendency of LLMs to generate rather brief responses which do not use the allowed maximal number of tokens. This unwillingness of models can be circumvented by prompting. Even so, the limited number of output tokens is typically too low and prevents effective extraction from long texts.
With a more sophisticated approach, the restriction (O) becomes irrelevant and only the restriction (I) will apply. The issue imposed by (O) is overcome by splitting the source document into smaller pieces which are extracted independently. A significant drawback is that the extracted information can be easily duplicated – extracted independently from multiple text pieces. Iterating the calls to the LLM with instruction to continue with already started extraction, i.e., continuing with the extraction in a single thread, helps to extract more information and avoid duplication. As we insist on continuation, more and more information is added and the extraction is more thorough, at least to some point – this will be addressed in detail in Section 5.1. Further, a lower number of duplicates is found due to the extraction history, i.e., all information extracted until present, which is kept within the thread.
@@ 5 Quality of extraction
Once the information is extracted from data into a structured form defined by the chosen schema, e.g., Figure 1, the quality of such extraction is important to evaluate. In practice, it is very rare to be equipped with ground truth and its human generation requires vast expertise in the scope of data and a ridiculous amount of time. Therefore we adopt methods from [4]. They examine semantic similarity, relevance, redundancy, and bias and compound these into a single score called SUSWIR, for details see Appendix A. The score and its subparts are very useful when comparing distinct extractions among themselves, e.g., we can use it to find an optimal number of iterated LLM calls. Unfortunately, the score does not represent an absolute way of evaluation. It does not provide a complete insight into the task – some information (= entities) can be missing, misclassified or their properties not filled in correctly. To come up with a robust and general solution we generalize the NIAH test, which is commonly used to measure the ability of LLMs to process long documents, cf. [6].
@@ 5.1 Iterated LLM calls
Since the first LLM extraction is typically not exhaustive, iterating the extraction process helps with the completeness of extraction. To improve the quality of extraction, we ask LLM to process the document again and search for other entities which were not extracted yet. A question arises: What is the optimal number of iterations? It is desirable to stop when additional LLM call will return no or only a few new entities. The answer however depends heavily on the text being extracted and on the chosen schema. Below, we present a small comparative study regarding the contribution of iterated extraction to its quality. We interpret the extracted structured data, e.g., Figure 3, as a summary of the original text document. To measure the quality of the summary we adopt the scores from [4] (a convex combination of these scores creates the overall SUSWIR metric), namely semantic similarity, relevance, and redundancy avoidance. We use a modified bias avoidance score from [4] and add two new scores, relevance spread, and incompleteness score. See Appendix A for more details.
Consider document which length is approximately 12k chars. Table 2 compares the content of the document with extracted information created iteratively by succeeding LLM calls. Each iteration enriches the extracted information, but the benefit decreases. From the third iteration, i.e., after four LLM calls, the majority of scores in Table 2 are either getting worse or stagnating (the arrows following the score name indicate the direction in which the score improves). It is obvious that shorter and longer text will require less or more iterations to extract majority of information without reducing its semantic and factually relevant meaning, respectively. Further, the risk that the LLM will suffer from hallucinations increases as we observe a growth of bias. In the rest of the paper we use three iterations to extract documents of approximate length 12k chars within all extractions (if not stated otherwise).
@@ 5.2 Test the quality
This section introduces a robust and versatile score to objectively measure the quality of IE. Assuming the structure is imposed by some schema, see Section 3.1, we would like to measure the IE quality as a portion of successfully extracted entities, i.e., the accuracy of name entity recognition (NER) task taking into account even the context captured by entity properties. Unfortunately, such an experiment is unfeasible without labeled data. As a consequence, it is unfeasible in many specialized tasks because of the absence of suitable labeled data unseen by LLM models. This can be the case with very recent datasets as well as proprietary datasets. To overcome this issue we use inspiration by NIAH test to build up an automatic and general procedure to access the quality of IE tasks.
@@ 5.2.1 Needles
A ‘needle’ in our context represents an entity. It is created according to the chosen schema, i.e., a list of types we want to extract from the document. We use an LLM to generate a short paragraph introducing a new original (not appearing in the document) entity, but still relevant to the scope of the document, for an example see Figure 2, and for more details on generation process see Appendix B2. This artificial paragraph, the needle, is then placed into the document body at random (taking into the account natural units within the text as sentences, paragraphs, etc. if applicable). Moreover, the needle is accompanied with several properties, namely we assign to the needle a name, short description and keywords, see Figure 2. This additional properties are assigned to the needle by the LLM.
@@
5.2.2 Multiple infused needle extraction accuracy
To measure the quality of extraction we propose a multiple infused needle extraction accuracy (MINEA) score. Its computation combines the approach of NIAH evaluation and NER task. We scatter several needles at random over the text document body (such that the inserted needles fill 10 to 30% of the enriched text) and measure how many of them were successfully extracted. Since we know what exactly was inserted, we know what should be extracted. Then we can objectively measure the quality of extraction on these new entities and moreover, we can compare extracted information from the document with and without needles. Table 3 shows extraction accuracy – MINEA score – total and per schema type – measured on a vast corpus of business documents with predefined schema consisting of types ‘BioChemEntity’, ‘Event’, ‘Insight’, ‘Legislation’, ‘MedicalCondition’, ‘OpportunityArea’, ‘Person’, ‘Product’, ‘Project’, ‘Substance’ and ‘Thing’.
@@ 5.2.3 Identification of needles
Matching the generated needles with extracted entities imposes a challenge and mostly depends on the formulation of needles. If the needles are too complex or too vague, the straightforward identification changes into a serious problem. For this reason, we equip the needles with additional properties which are then used to compare the needles with extracted entities and to decide whether the needles were extracted successfully or not.
@@
    an entity with some number of keywords perfectly matching the needle keywords is found, the number is determined by the threshold parameter determining the percentage of keywords to be matched;
@@
Note that other conditions can be constructed, e.g., based on the short description instead of keywords, etc. Table 4 shows whether the conditions are fulfilled in the example illustrated by Figures 2 and 3. Namely, the condition n is not satisfied (‘AI Clan Meeting’ $\neq$ ‘AI Meeting’, ‘Graph Index’ $\neq$ ‘GRIX’). Condition ns is satisfied only for needle representing an entity of type ‘Event’ (‘AI Clan Meeting’ can be found in the extracted information). There are three keywords out of the six assigned to the needle representing the entity of type ‘Event’ which match the keywords of an extracted entity, hence k0.5 is, and k0.6, k0.7 are not satisfied (there is an entity within the extracted information with $50\%$ of keywords being the same as the keywords of the needle). In the case of the second needle, there are four such keywords, therefore k0.5 and k0.6 are satisfied. Finally, both needles are identified within the extracted information by an LLM.
Table 5 shows scores (ratios of successfully extracted entities) based on the above criteria in the case of our business documents. The types of inserted needles are ’BioChemEntity’, ’Country’, ’Event’, ’Insight’, ’Legislation’, ’Person’, ’Product’, ’Project’ and ’Substance’. Matching the needle and entity name usually does not perform well if the name is prone to modification (e.g., person name with and without title), or if the entity is easy to be misclassified (an entity of type ‘Country’ was often extracted as ‘Place’ whose name did not match the country name). Searching for a needle name in all extracted information gives very accurate results if the entities are well characterized by their name (compare for example types ‘Person’ and ’Legislation’ with type ’Insight’ where the name is not a natural attribute). Matching the needle and entity keywords depends on the threshold parameter – with a lower proportion of keywords that have to match the score value increases and the reliability of the entity identification decreases. An LLM performs well the entity identification and it is an important criterion in the case of more creative types such as ‘Insight’. Finally, the MINEA score for each type is taken as the maximum of the scores (the values are highlighted).
@@
5.2.4 Model comparison
MINEA score can be used to compare the performance of distinct LLMs, see Table 6. A corpus of documents is infused by needles representing entities whose types match the schema introduced in Section 5.2.2. Three OpenAI LLMs [^4] https://platform.openai.com/docs/models are used to extract a relevant information under the same setting (the same model parameters such as temperature, the same number of iterations, the same prompting, etc.). Model gpt-3.5-turbo is outperformed by gpt-4-turbo by almost 15% and gpt-4-turbo is outperformed by gpt-4o model by another 12%. Note that the achieved accuracy is lower than presented in Table 3, since only one iteration instead of three was performed in order to reduce the computational time.
Table 6: LLMs comparison using MINEA score.
model gpt-3.5-turbo gpt-4-turbo gpt-4o
MINEA 0.449198 0.593583 0.716578